OpenStudy (anonymous):

What is the slope of (4.5,15.5) , (5,16)

OpenStudy (anonymous):

(y2-y1)/(x2-x1) =(16-15.5)/(5-4.5) =.5/.5 =1

OpenStudy (anonymous):

How would you figure out the y-intercept?

OpenStudy (anonymous):

first, you have to form and equation... y-y1 = m ( x- x1) y-16 = 1( x-5) y=x-5+16 y= x+11--------------1 now, comparing this equation with slope intercept form of equation, i.e y=mx+c, c = 11 i.e, y-intercept = 11 or, you can do in another way too, plug in value of x =0 , and you will get y-intercept! :)

OpenStudy (anonymous):

So when they ask for the slope intercept form, is that an equation?

OpenStudy (anonymous):

yes the one I have marked as----- 1

OpenStudy (anonymous):

That is what I thought. Thank you very much(:

OpenStudy (anonymous):

welcome